OpenStudy (anonymous):

The units digit of a perfect square is 6. What are the possible values of the tens digit?

OpenStudy (freckles):

I would play around with the numbers... like we know 4^2=16 and 6^2=36 what is 14^2, 16^2, 24^2,... what are the tens digits in each of those ...

OpenStudy (anonymous):

x^2=6 mod 10 thus x=4,6 mod 10 thus x=4+10k or 6+10k thus x^2=(4+10k)^2 or (6+10k)^2 find x^2 mod 100